OpenStudy (anonymous):

Use a infinite sum to estimate the average value of the function on the given interval by partioning the interval and evaluating the function at the midpoints of the subintervals. f(x)=x^2-3 on [-3, 7]divided into 5 subintervals

OpenStudy (anonymous):

Since 5 is not infinity, it wouldn't be an infinite sum.

OpenStudy (anonymous):

Interval size is \[ \frac{b-a}{n}=\frac{7-(-3)}{5} = 2 \]

OpenStudy (anonymous):

Our values for \(x\), will be \[ x_i^* = \Delta x(i-1)+\Delta{x}/2 = 2(i-1)+1 = 2i-1 \]

OpenStudy (anonymous):

All together: \[ \sum_{i=1}^{5}f(x_i^*)\Delta x = \sum_{i=1}^{5}f(2i-1)\cdot2 = \sum_{i=1}^{5}((2i-1)^2-3)\cdot 2 \]

OpenStudy (aum):

\[ x_i^* = -3 + \Delta x(i-1)+\Delta{x}/2 = -3 + 2(i-1) + 1 = -3 + 2i - 2 + 1 = 2i - 4 \]